# Automatic and Structure-Aware Sparsification of Hybrid Neural ODEs

### v1

@@ Acknowledgment
A1.2 Proof for Group LASSO Equivalence
@@
and ($\text{and }~{}\sum_{(u,v)\in E^{a,c}}\left(\lambda_{1}|w_{(u,v)}|+\lambda_{2}\|\Theta_{(u,v)}\|_{2}^{2}\right)+\lambda_{2}\|\tilde{\Theta}\|_{2}^{2},$
@@
$+\lambda_{2}\|\tilde{\Theta}\|_{2}^{2}+\sum_{(u,v)\in E^{a,c}}\left(\lambda_{1}|w_{(u,v)}|+\lambda_{2}\frac{\|\Gamma_{(u,v)}\|_{2}^{2}}{w_{(u,v)}^{2}}\right).$ (4)
@@
$w^{*}_{(u,v)}=\left(\frac{2\lambda_{2}\|\Gamma_{(u,v)}\|_{2}}{\lambda_{1}}\right)^{\frac{1}{3}}.$
@@
$+\lambda_{2}\|\tilde{\Theta}\|_{2}^{2}+2\sqrt{\lambda_{1}\lambda_{2}}\sum_{(u,v)\in E^{a,c}_{M}}\|\Gamma_{(v,u)}\|_{2}$
@@
A2.2 Features, Units and Interpolation
@@
A2.3 Interpolating Basal Flow Rate
@@
A2.4 Interpolating Bolus Insulin Rate
@@
A2.5 Interpolating Carbohydrate Intake Rate
@@
A2.6 Interpolating Heart Rate and Step Count
@@
A2.7 Choice of Time Grid and Discretization
@@
[…]
@@
A3.1 Synthetic Mechanistic Graph
@@
For all experiments, we use the Adam optimizer [Kingma and Ba, 2015] to perform stochastic gradient descent. We initialize all model weights with the PyTorch default setting with seed $2024+{r},$ where $r$ increases by 1 starting from 0 for each experiment repetition (40 repeats for synthetic and 10 repeats for real-world experiments) unless state otherwise in the detailed model computation algorithm below.
@@
A3.3 MNODE without Reduction (MNODE_NR)
@@
The LSTM has $2$ layers and $|V_{\text{uva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in the directed graph) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 00 and activation ReLu. We place $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A3.4 MNODE reduced by domain knowledge (MNODE_DK)
We take the reduced UVA-Padova model from Appendix C of Zou et al. [2024], and use its directed graph representation denoted as $G_{\text{ruva}}$ to construct MNODE_DKR. Other than the graph used by MNODE, model computation is exactly the same the MNODE_NR.
@@
The LSTM has $2$ layers and $|V_{\text{ruva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in the reduced graph) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 00 and activation ReLu. We place $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A3.5 MNODE with Hybrid Graph Sparsification (MNODE_HGS)
@@
The LSTM has $2$ layers and $|V^{a}_{\text{uva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in $G^{a,c}_{\text{uva}}$) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 00 and activation ReLu. We place $L_{1}$ regularization on all the edge weights with penalty hyper-parameter $\lambda_{1}$, $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A3.6 MNODE reduced by exclusive group LASSO (MNODE_EGL)
@@
A3.7 MNODE reduced by elastic net (MNODE_EN)
@@
A3.8 MNODE reduced by Neural Sparse (MNODE_NS)
@@
A3.9 MNODE reduced by greedy search (MNODE_GD)
@@
  MinLoss$=10^{7}$
  while Stop $=$ FALSE do
@@
A3.10 MNODE reduced by random search (MNODE_RD)
The random search randomly picks 5 sub-graphs that has contains $1-p$ percent of edges and select the best one. Specifically:
@@
[…]
@@
[…]
@@
[…]
     for $i=1:q-1$ do
[…]
@@
[…]
@@
A4.2 Real World Experiment
@@
so that we can estimate bias and variance separately. Under this assumption, noise is zero and $\text{EPE}=\text{bias}^{2}+\text{variance}.$ Specifically, in this paper we use a modified K-fold cross-validation to construct unbiased estimators for variance and bias. First, split $D$ into $K$ disjoint subsets of equal size $D=\bigsqcup_{k=1}^{K}D_{k}$, denote $D^{(-k)}=D\setminus D_{k}$. Then, each $D^{(-k)}$ consists of $K-1$ disjoint subsets $D_{i}:$
@@
A5.1 Summary Diagram
At a high level, UVA-Padova can be summarized by the diagram in Figure 6, which is taken from Figure 1 in Man et al. [2014]. It divides the complex physiological system into 10 subsystems, which are linked by key causal states such as Rate of Appearance, Endogenous Glucose Production and Utilization. Next, we will introduce each subsystem one by one and also explain the physiological meanings behind state variables.
Figure 6: UVA/Padova Simulator S2013, taken from Figure 1 of Man et al. [2014]
A5.2 Glucose Subsystem
@@
A5.4 Glucose Rate of Appearance
@@
A5.5 Endogenous Glucose Production
@@
$risk=\begin{cases}0&G_{b}\leq G\\ 10(\log(G)-\log(G_{b}))^{2r_{2}}&G_{th}\leq G<G_{b}\\ 10(\log(G_{th})-\log(G_{b}))^{2r_{2}}&G<G_{th}\\ \end{cases}$ (24)
$F_{cns}$: Glucose Independent Utilization Constant, $X$: Insulin Action on Glucose Utilization, $I_{b}$ Basal Insulin Concentration Constant, $risk$ Hypoglycemia Risk Factor, $G_{b}$ Basal Glucose Concentration Parameter, $G_{th}$ Hypoglycemia Glucose Concentration Threshold.
A5.7 Renal Excretion
@@
A5.8 Subcutaneous Insulin Kinetics
@@
A5.9 Subcutaneous Glucose Kinetics
@@
A5.10 Glucagon Secretion and Kinetics
@@
A5.11 Subcutaneous Glucagon Kinetics
@@
A6.1 Additional metrics for synthetic experiments
@@
A6.2 Ablation study on HGS under limited data, true sparsity regime with comprehensive starting graph

### v2

@@ Appendix A Appendix
A2.2 Proof for Group LASSO Equivalence
@@
$\text{and }~\sum_{(u,v)\in E^{a,c}}\left(\lambda_{1}|w_{(u,v)}|+\lambda_{2}\|\Theta_{(u,v)}\|_{2}^{2}\right)+\lambda_{2}\|\tilde{\Theta}\|_{2}^{2},$
@@
$\lambda_{2}\|\tilde{\Theta}\|_{2}^{2}+\sum_{(u,v)\in E^{a,c}}\left(\lambda_{1}|w_{(u,v)}|+\lambda_{2}\frac{\|\Gamma_{(u,v)}\|_{2}^{2}}{w_{(u,v)}^{2}}\right).$ (4)
@@
$w^{*}_{(u,v)}=\left(\frac{2\lambda_{2}\|\Gamma_{(u,v)}\|^{2}_{2}}{\lambda_{1}}\right)^{\frac{1}{3}}.$
@@
$\lambda_{2}\|\tilde{\Theta}\|_{2}^{2}+2\sqrt{\lambda_{1}\lambda_{2}}\sum_{(u,v)\in E^{a,c}_{M}}\|\Gamma_{(v,u)}\|_{2}$
@@
A3.2 Features, Units and Interpolation
@@
A3.3 Interpolating Basal Flow Rate
@@
A3.4 Interpolating Bolus Insulin Rate
@@
A3.5 Interpolating Carbohydrate Intake Rate
@@
A3.6 Interpolating Heart Rate and Step Count
@@
A3.7 Choice of Time Grid and Discretization
@@
Average Carb Intake Rate $M_{i}=\frac{1}{\Delta t}\int_{t_{i}}^{t_{i+1}}f_{m}(t)\,dt$
@@
A4.1 Synthetic Mechanistic Graph
@@
For all experiments, we use the Adam optimizer (Kingma & Ba, 2015) to perform stochastic gradient descent. We initialize all model weights with the PyTorch default setting with seed 2024+r,2024+{r}, where $r$ increases by 1 starting from 0 for each experiment repetition (40 repeats for synthetic and 10 repeats for real-world experiments) unless state otherwise in the detailed model computation algorithm below.
@@
A4.3 MNODE without Reduction (MNODE_NR)
@@
The LSTM has $2$ layers and $|V_{\text{uva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in the directed graph) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 0 and activation ReLu. We place $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A4.4 MNODE reduced by domain knowledge (MNODE_DK)
We take the reduced UVA-Padova model from Appendix C of Zou et al. (2024), and use its directed graph representation denoted as $G_{\text{ruva}}$ to construct MNODE_DKR. Other than the graph used by MNODE, model computation is exactly the same the MNODE_NR.
@@
The LSTM has $2$ layers and $|V_{\text{ruva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in the reduced graph) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 0 and activation ReLu. We place $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A4.5 MNODE with Hybrid Graph Sparsification (MNODE_HGS)
@@
The LSTM has $2$ layers and $|V^{a}_{\text{uva}}|$ hidden dimension (i.e. this corresponds to the number of nodes in $G^{a,c}_{\text{uva}}$) and for the same reason we set the 1st features of the initial condition state vector to be observed initial glucose level $s_{1}^{t_{0}}$. All MLPs have $2$ hidden layers and $16$ hidden units with dropout 0 and activation ReLu. We place $L_{1}$ regularization on all the edge weights with penalty hyper-parameter $\lambda_{1}$, $L_{2}$ regularization on all MLP parameters with penalty hyper-parameter $\lambda_{2}$, and train the model with learning rate $lr$. We tune these hyper-pameters with grid search on the following grid:
@@
A4.6 MNODE reduced by exclusive group LASSO (MNODE_EGL)
@@
A4.7 MNODE reduced by elastic net (MNODE_EN)
@@
A4.8 MNODE reduced by Neural Sparse (MNODE_NS)
@@
A4.9 MNODE reduced by greedy search (MNODE_GD)
@@
[…]
while Stop == FALSE do
@@
A4.10 MNODE reduced by random search (MNODE_RD)
The random search randomly picks 5 sub-graphs that has contains 1p1-p percent of edges and select the best one. Specifically:
@@
$\text{encoder\_in}=\text{concatenate}(D^{\text{P}},X^{\prime},\text{dim}=\text{time})$ (concatenating all inputs to form a masked context)
@@
  $\text{decoder\_in}=\text{concatenate}(s_{1}^{t_{0}},s_{1}^{t_{1:q-1}})$ (expected output shifted to the right)
  $\text{decode\_out}=\text{Transformer}(\text{encoder\_in},\text{decoder\_in},\text{decoder\_causal\_mask})$
@@
  $\text{decoder\_in}=\text{concatenate}(s_{1}^{t_{0}},\mathbf{0}\in\mathbb{R}^{q-1})$
  for i=1:q1i=1:q-1 do
   $\text{decode\_out}=\text{Transformer}(\text{encoder\_in},\text{decoder\_in})$
   $\text{decoder\_in}_{i+1}=\text{decode\_out}_{i}$
@@
  $\text{decode\_out}=\text{Transformer}(\text{encoder\_in},\text{decoder\_in})$
@@
A5.2 Real World Experiment
@@
so that we can estimate bias and variance separately. Under this assumption, noise is zero and $\text{EPE}=\text{bias}^{2}+\text{variance}.$ Specifically, in this paper we use a modified K-fold cross-validation to construct unbiased estimators for variance and bias. First, split $D$ into $K$ disjoint subsets of equal size $D=\bigsqcup_{k=1}^{K}D_{k}$, denote $D^{(-k)}=D\setminus D_{k}$. Then, each $D^{(-k)}$ consists of K1K-1 disjoint subsets $D_{i}:$
@@
A6.1 Summary Diagram
At a high level, UVA-Padova can be summarized by the diagram in Figure 6, which is taken from Figure 1 in Man et al. (2014). It divides the complex physiological system into 10 subsystems, which are linked by key causal states such as Rate of Appearance, Endogenous Glucose Production and Utilization. Next, we will introduce each subsystem one by one and also explain the physiological meanings behind state variables.
Figure 6: UVA/Padova Simulator S2013, taken from Figure 1 of Man et al. (2014)
A6.2 Glucose Subsystem
@@
A6.4 Glucose Rate of Appearance
@@
A6.5 Endogenous Glucose Production
@@
[…]
$F_{cns}$: Glucose Independent Utilization Constant, $X$: Insulin Action on Glucose Utilization, $I_{b}$ Basal Insulin Concentration Constant, riskrisk Hypoglycemia Risk Factor, $G_{b}$ Basal Glucose Concentration Parameter, $G_{th}$ Hypoglycemia Glucose Concentration Threshold.
A6.7 Renal Excretion
@@
A6.8 Subcutaneous Insulin Kinetics
@@
A6.9 Subcutaneous Glucose Kinetics
@@
A6.10 Glucagon Secretion and Kinetics
@@
A6.11 Subcutaneous Glucagon Kinetics
@@
A7.1 Additional metrics for synthetic experiments
@@
A7.2 Ablation study on HGS under limited data, true sparsity regime with comprehensive starting graph
@@
Appendix A8 Instability of Dynamical Systems with Cycles: A toy example
Here we discuss how cycles/loops in dynamical systems can lead to instability. There are three sources of numerical stability: blowing-up, exploding gradient and stiffness.
A8.1 Blowing-up
Consider a simple 2 state system with a circular dependence:
$\frac{ds_{1}(t)}{dt}=as_{1}(t)+bs_{2}(t)$
$\frac{ds_{2}(t)}{dt}=cs_{1}(t)+ds_{2}(t)$
The system will blow up when the Jacobian of the system,
$J=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$,
has eigenvalues with positive real parts. That means that if one solves the system using the forward Euler method with a step size $h$, then the system with the update rule
$\left(\begin{array}[]{c}s_{1}(t+kh)\\ s_{2}(t+kh)\end{array}\right)=\left(hJ+I\right)^{k}\left(\begin{array}[]{c}s_{1}(t)\\ s_{2}(t)\end{array}\right)$
would blow up as $k$ increases. Let us then consider the two eigenvalues of $J$:
$\lambda_{\pm}(J)=\frac{1}{2}\left(a+d\pm\sqrt{(a-d)^{2}+4bc}\right).$
If $bc>\max(0,ad)$, then $\lambda_{+}(J)$ is a positive number making the system unbounded, even if $a$ and $d$ are negative.
On the other hand, suppose the system is acyclic except for self-loops (exactly what our step 1 is doing), say $ds_{1}(t)/dt$ depends on $s_{2}(t)$ but $ds_{2}(t)/dt$ does not depend on $s_{1}(t)$ anymore, then $J$ becomes
$J=\begin{bmatrix}a&b\\ 0&d\end{bmatrix}$
with $a$ and $d$ as its eigenvalues. Then, the system is allowed to freely model how $s_{2}(t)$ affects $ds_{1}(t)/dt$ without concerns of explosion (no constraint on $b$), as long as $a$ and $d$ are negative.
A8.2 Exploding gradient
Now suppose we discretize the system and solve it with RNNs, then the gradient will also behave like $\left(J+I\right)^{k},$ which will cause the RNNs to have exploding gradients when $k$ is large.
A8.3 Stiffness
It is also well-known in both the physics and the neural ODE community that stiffness of the system can also cause numerical instability (Kim et al., 2021; Worsham & Kalita, 2025).
Consider the dissipative version of the 2-state system with a circular dependence:
$\frac{ds_{1}(t)}{dt}=-s_{1}(t)+bs_{2}(t)$
$\frac{ds_{2}(t)}{dt}=cs_{1}(t)-s_{2}(t)$
The eigenvalues of the system are
$\lambda_{\pm}(J)=-1\pm\sqrt{bc}.$
If $0<bc<1,$ then $\lambda_{\pm}<0$ and the system blowing-up/exploding gradient are under control. However, the system can still suffer from stiffness, which is often defined as the ratio between the magnitude of the fastest to slowest stable rates:
$\kappa\equiv\frac{|\lambda_{-}|}{|\lambda_{+}|}=\frac{1+\sqrt{bc}}{1-\sqrt{bc}}.$
Therefore, the system stiffness can still blow up, if $bc\uparrow 1.$ On the other hand, without circular dependence, both eigenvalues are 1-1 and the stiffness is always 1 regardless of the value of $b$.
In summary, to ensure the stability of the ODE system, more complex constraints on model parameters are needed for a system with cycles than for a system without cycles.
Appendix A9 Tabulated Results
A9.1 Synthetic data experiments
We attach the tabulated version of experiment results on synthetic data in Table 4 to Table 9.
Table 4: Black-box comparison (mean $\pm$ standard error over 40 trials) (Sample Size $N=100$)
Model RMSE Peak RMSE ENP
True Sparsity
MNODE $\mathbf{0.1039\pm 0.0013}$ $\mathbf{0.4124}$ 757.55
BNODE $0.1237\pm 0.0021$ 1.2451 3686.78
S4D $0.1130\pm 0.0012$ 0.8247 $\mathbf{452.78}$
TCN $0.1214\pm 0.0049$ 1.8787 13671.23
Transformer $0.1699\pm 0.0039$ 1.2015 10892.65
LSTM $0.1191\pm 0.0021$ 0.9048 8510.78
Quasi Sparsity
MNODE $\mathbf{0.1056\pm 0.0014}$ $\mathbf{0.4971}$ 760.73
BNODE $0.1249\pm 0.0021$ 0.7979 3441.73
S4D $0.1137\pm 0.0012$ 0.6873 $\mathbf{456.38}$
TCN $0.1274\pm 0.0056$ 2.4566 13154.93
Transformer $0.1698\pm 0.0039$ 1.1801 10929.17
LSTM $0.1194\pm 0.0021$ 0.8659 8338.50
Table 5: Black-box comparison (mean $\pm$ standard error over 40 trials) (Sample Size $N=1000$)
Model RMSE Peak RMSE ENP
True Sparsity
MNODE $0.0981\pm 0.0010$ $\mathbf{0.2369}$ 730.25
BNODE $0.1009\pm 0.0011$ 0.5123 2219.60
S4D $0.0986\pm 0.0010$ 0.5624 $\mathbf{602.63}$
TCN $\mathbf{0.0875\pm 0.0010}$ 0.7938 14360.29
Transformer $0.2437\pm 0.0066$ 1.2028 10696.88
LSTM $0.1003\pm 0.0011$ 0.7060 8982.25
Quasi Sparsity
MNODE $0.0987\pm 0.0010$ $\mathbf{0.2377}$ 727.25
BNODE $0.1014\pm 0.0011$ 0.5887 2330.20
S4D $0.0991\pm 0.0010$ 0.4444 $\mathbf{647.00}$
TCN $\mathbf{0.0896\pm 0.0011}$ 0.6814 14173.83
Transformer $0.1420\pm 0.0027$ 0.8703 10576.15
LSTM $0.1008\pm 0.0011$ 0.6457 8981.73
Table 6: Reduction method comparison (mean $\pm$ standard error over 40 trials)
(Sample Size $N=100$) Refined Initial Graph.
Model RMSE Peak RMSE ENP
GL $\mathbf{0.1039\pm 0.0013}$ $\mathbf{0.3124}$ $\mathbf{757.55}$
EGL $0.1225\pm 0.0021$ 0.4826 1771.40
NS $0.1070\pm 0.0014$ 0.3833 1046.95
EN $0.1225\pm 0.0021$ 0.4820 1770.38
RD $0.1098\pm 0.0016$ 0.6919 1208.05
GD $0.1114\pm 0.0016$ 0.7097 1425.48
NR $0.1233\pm 0.0020$ 0.6476 1643.83
Table 7: Reduction method comparison (mean $\pm$ standard error over 40 trials)
(Sample Size $N=1000$) Refined Initial Graph.
Model RMSE Peak RMSE ENP
GL $\mathbf{0.0987\pm 0.0010}$ $\mathbf{0.2377}$ $\mathbf{727.25}$
EGL $0.1058\pm 0.0013$ 0.3055 1772.03
NS $0.1045\pm 0.0012$ 0.3492 1034.40
EN $0.1066\pm 0.0013$ 0.3101 1771.90
RD $0.1038\pm 0.0012$ 0.3423 1154.30
GD $0.1040\pm 0.0011$ 0.5046 1420.48
NR $0.1078\pm 0.0013$ 0.3805 1641.48
Table 8: Reduction method comparison (mean $\pm$ standard error over 40 trials)
(Sample Size $N=100$) Comprehensive graph.
Model RMSE Peak RMSE ENP
GL $\mathbf{0.1040\pm 0.0013}$ $\mathbf{0.3892}$ $\mathbf{1468.25}$
EGL $0.1415\pm 0.0031$ 0.8537 5098.98
NS $0.1845\pm 0.0039$ 1.1209 3565.75
EN $0.1794\pm 0.0050$ 0.9047 5094.78
RD $0.1236\pm 0.0022$ 0.5097 3905.25
GD $0.3129\pm 0.0063$ 1.8013 5053.78
NR $0.1235\pm 0.0020$ 0.4509 4933.95
Table 9: Reduction method comparison (mean $\pm$ standard error over 400 trials)
(Sample Size $N=1000$) Comprehensive graph.
Model RMSE Peak RMSE ENP
GL $\mathbf{0.0992\pm 0.0010}$ $\mathbf{0.2339}$ $\mathbf{1415.03}$
EGL $0.1287\pm 0.0022$ 0.5938 5097.38
NS $0.1758\pm 0.0035$ 0.6184 3881.53
EN $0.1287\pm 0.0022$ 0.5397 5095.23
RD $0.1278\pm 0.0021$ 0.4183 4410.73
GD $0.3415\pm 0.0058$ 1.5982 5048.73
NR $0.1165\pm 0.0017$ 0.3507 4946.93
A9.2 Real-world data experiments
We attach the tabulated version of experiment results on real-world data in Table 10 and Table 11
Table 10: Predictive performance (mean $\pm$ standard error over 10 trials).
Model RMSE MAPE Corr Acc.
MNODE_NR $36.19\pm 0.33$ $0.230\pm 0.002$ $0.649\pm 0.006$ $0.760\pm 0.004$
MNODE_DK $36.58\pm 0.60$ $0.229\pm 0.003$ $0.657\pm 0.006$ $0.765\pm 0.004$
MNODE_HGS12 $35.92\pm 0.31$ $0.227\pm 0.002$ $0.657\pm 0.004$ $0.768\pm 0.002$
MNODE_HGS1 $36.12\pm 0.40$ $0.229\pm 0.003$ $0.648\pm 0.005$ $0.764\pm 0.005$
MNODE_HGS2 $35.96\pm 0.30$ $0.229\pm 0.002$ $0.650\pm 0.005$ $0.766\pm 0.002$
MNODE_HGS $\mathbf{35.22}\pm 0.25$ $\mathbf{0.223}\pm 0.002$ $\mathbf{0.682}\pm 0.003$ $\mathbf{0.786}\pm 0.002$
MNODE_HGS3 $36.13\pm 0.41$ $0.229\pm 0.002$ $0.651\pm 0.005$ $0.768\pm 0.003$
MNODE_HGS13 $35.82\pm 0.37$ $0.227\pm 0.002$ $0.654\pm 0.007$ $0.773\pm 0.004$
MNODE_HGS23 $35.95\pm 0.27$ $0.228\pm 0.002$ $0.669\pm 0.003$ $0.769\pm 0.003$
MNODE_EGL $35.86\pm 0.25$ $0.227\pm 0.001$ $0.650\pm 0.005$ $0.760\pm 0.003$
MNODE_EN $35.93\pm 0.29$ $0.227\pm 0.002$ $0.651\pm 0.005$ $0.760\pm 0.002$
MNODE_NS $35.55\pm 0.28$ $0.227\pm 0.002$ $0.655\pm 0.005$ $0.768\pm 0.003$
MNODE_GD $36.70\pm 0.44$ $0.229\pm 0.002$ $0.655\pm 0.007$ $0.765\pm 0.003$
MNODE_RD $35.78\pm 0.41$ $0.227\pm 0.003$ $0.662\pm 0.002$ $0.769\pm 0.003$
BNODE $37.08\pm 0.25$ $0.260\pm 0.002$ $0.666\pm 0.003$ $0.759\pm 0.003$
S4D $42.91\pm 0.39$ $0.283\pm 0.003$ $0.629\pm 0.005$ $0.724\pm 0.002$
LSTM $40.69\pm 0.39$ $0.266\pm 0.003$ $0.666\pm 0.003$ $0.733\pm 0.003$
TCN $41.09\pm 0.44$ $0.277\pm 0.003$ $0.672\pm 0.003$ $0.725\pm 0.002$
Transformer $46.29\pm 0.51$ $0.283\pm 0.003$ $0.592\pm 0.004$ $0.664\pm 0.009$
Table 11: Complexity and peak-error metrics. Variance shows mean $\pm$ standard error over 10 trials; ENP and peak metrics are means only.
Model Variance ENP Peak RMSE Peak MAPE
MNODE_NR $125.8\pm 17.9$ $10684$ $189.6$ $1.401$
MNODE_DK $149.2\pm 34.3$ $6956$ $183.3$ $1.464$
MNODE_HGS12 $119.9\pm 17.1$ $9033$ $202.7$ $1.469$
MNODE_HGS1 $136.1\pm 21.5$ $8848$ $177.9$ $1.398$
MNODE_HGS2 $115.4\pm 20.5$ $10643$ $193.7$ $1.426$
MNODE_HGS $\mathbf{76.4}\pm 17.9$ $\mathbf{7551}$ $\mathbf{123.4}$ $\mathbf{1.222}$
MNODE_HGS3 $101.5\pm 23.8$ $7966$ $167.8$ $1.313$
MNODE_HGS13 $96.6\pm 17.2$ $7735$ $169.3$ $1.354$
MNODE_HGS23 $96.4\pm 19.9$ $8054$ $166.6$ $1.354$
MNODE_EGL $92.0\pm 12.4$ $8326$ $169.4$ $1.321$
MNODE_EN $101.7\pm 19.0$ $8548$ $169.0$ $1.239$
MNODE_NS $96.8\pm 21.5$ $8730$ $161.3$ $1.330$
MNODE_GD $124.7\pm 23.8$ $8861$ $260.6$ $1.349$
MNODE_RD $107.0\pm 22.4$ $8955$ $184.8$ $1.322$
BNODE $332.5\pm 17.4$ $8596$ $190.9$ $1.541$
S4D $291.4\pm 11.3$ $8099$ $194.3$ $1.667$
LSTM $317.8\pm 27.9$ $8102$ $178.6$ $1.437$
TCN $384.5\pm 23.4$ $8261$ $161.6$ $1.803$
Transformer $509.0\pm 27.8$ $8122$ $210.2$ $1.953$